# Monadic Second Order Logic of the Real Order with the Set Quantifiers ranging over the Borel Sets

### v4

@@
I thank Joan Bagaria for the invitation to visit him at the University of Barcelona in April 2025. The conversation we had gave me a crucial insight. Namely, I showed Joan the idea of well-levelled orders and he proposed that this could probably be used in descriptive set theory. I went to work on this and connected it with Rabin’s work and solving the problem of MSO on the Borel sets, about which I had been thinking about on and off for a while. The idea of writing a joint paper on this did not converge, due to Joan’s other projects. I thank Andreas Blass, Jouko Vддnдnen and Andrйs Villaveces on very useful comments on the preprint of the paper.
@@ 2. Well-levelled orders and their basic properties
In the forward direction, let us write $P$ as the union of $\sim$-equivalence classes, which we shall enumerate as $\{C_{\alpha}:\alpha<\kappa\}$. Then each $C_{\alpha}$ is a linear order, since $p\sim q$ implies that $p$ and $q$ are comparable. Let $\alpha<\kappa$ and let $p\in C_{\alpha}$. We have that $p\sim q$ with some $q$ which is $P$-minimal, by the fact that ${\mathsf{wl}}(P)=0$. Hence $q\in C_{\alpha}$ and since $C_{\alpha}$ is linear, $q$ is also the unique minimal element of $C_{\alpha}$. Denote such $q$ by $q_{\alpha}$. Then we conclude as in (1) that $C_{\alpha}$ is the well-order $C_{\alpha}=\bigcup_{p\in C_{\alpha}}[q_{\alpha},p]$. We shall show that there is no relation between the elements of $C_{\alpha}$ and $C_{\beta}$ for $\alpha\neq\beta$.
@@
A path in $T^{\alpha}$ is a linearly ordered subset set with no upper bound. Note that by the fact that $T^{\alpha}$ is well-founded, every path is order isomorphic to an ordinal. An $\omega$-path is a path that is order isomorphic to $\omega$. A branch is a maximal path. The set of all branches in $T^{\alpha}$ is denoted by $T^{\alpha}]$.
@@ 2.1. The Polish Topology of $T^{1+\alpha}]$
Suppose that $(X_{n},\rho_{n})_{n<\omega}$ is a sequence of non-empty complete metric spaces with metrics bounded by $1$. Then the product $\prod_{n<\omega}X_{n}$ admits a complete metric given by
$\rho(f,g)=\sum_{n<\omega}2^{-(n+1)}\rho_{n}(f(n),g(n)).$
@@
Theorem 2.14.
For every $\alpha<\omega_{1}$, there is a canonical injection $e_{\alpha}$ from $T^{1+\alpha}$ to $\prod_{1+\alpha}T$,which lifts to an injection $F_{\alpha}$ from $T^{1+\alpha}]$ to $\prod_{1+\alpha}[T]$ given by
$F_{\alpha}(\rho)=\{e_{\alpha}(t):\,t\in\rho\}.$
The image of $F_{\alpha}$ is a closed subset of $\prod_{1+\alpha}[T]$.
@@
The proof is by induction on $\alpha$. A part of the inductive assumption is that for every $\beta<\alpha$ and $t\in T^{1+\alpha}$ we have $e_{\alpha}(t)\upharpoonright[1+\beta)=e_{\beta}(t\upharpoonright[1+\beta))$.
For $\alpha=0$ the required injection $e_{0}$ is the identity and it lifts to $F_{0}$ which is also the identity. The induction hypothesis is vacuous.
For $\alpha+1$, we first define $e_{\alpha+1}$, given by $e_{\alpha+1}(t)=(t(\alpha),e_{\alpha}(t\upharpoonright[1,\alpha)))$. Since we have assumed that $e_{\alpha}$ is an injection, then clearly so is $e_{\alpha+1}$. Note that the induction hypothesis is preserved by this definition.
Suppose that $\rho$ is a branch of $T^{1+\alpha+1}$. We have defined
$F_{\alpha+1}(\rho)=\{e_{\alpha+1}(t):\,t\in\rho\}=\{(t(\alpha),e_{\alpha}(t\upharpoonright[1+\alpha))):\,t\in\rho\}.$
Since $\rho^{\prime}=\rho\upharpoonright[1+\alpha)$ is a branch of $T^{1+\alpha}$, we have by the inductive hypothesis that $F_{\alpha}(\rho^{\prime})$ is a branch of $\prod_{1+\alpha}T$. Moreover, we have that $F_{\alpha+1}(\rho)\upharpoonright[1+\alpha)=F_{\alpha}(\rho^{\prime})$. On the other hand, $F_{\alpha+1}(\rho)(\alpha)=\rho(\alpha)$, where by this notation we mean the projection on the coordinate $\alpha$. Since $\rho$ is a branch of $T^{1+\alpha+1}$, we have that $\rho(\alpha)$ is a branch of $T$ and hence we conclude that $F_{\alpha+1}(\rho)$ is a branch of $\prod_{1+\alpha+1}T$.
The image of $F_{\alpha+1}$ is closed in $\prod_{1+\alpha}[T]$ because its complement is $\prod_{1+\alpha}[T]\setminus[{\rm Im}(F_{\alpha})\times[T]]$, which is open in $\prod_{1+\alpha+1}[T]$ by the inductive hypothesis and the definition of the product topology.
For $\alpha>0$ a limit, given $t\in T^{\alpha}$ we let $e_{\alpha}(t)$ be the element of $\prod_{\alpha}T$ such that for every $\beta<\alpha$ we have $e_{\beta}(t\upharpoonright[1+\beta))=e_{\alpha}(t\upharpoonright[1\upharpoonright\beta))$. This is well defined by the induction hypothesis and it conserves the induction hypothesis at this step.
Suppose that $\rho$ is a branch of $T^{\alpha}$. By the inductive hypothesis we have that for every $\beta<\alpha$ the set $F_{\beta}(\rho\upharpoonright[1+\beta))$ is an element of $\prod_{1+\beta}[T]$, so $F_{\alpha}(\rho)$ is indeed an element of $\prod_{\alpha}[T]$. Suppose that $f\in\prod_{\alpha}[T]$ is not in the image of $F_{\alpha}$. This implies that there is $\beta<\alpha$ such that $f\upharpoonright[1+\beta)$ is not in the image of $F_{\beta}$. In particular there is a basic open set $O$ of $\prod_{1+\beta}[T]$ containing $f\upharpoonright[1+\beta)$ and disjoint from the image of $F_{\beta}$. The inverse projection of $O$ in $\prod_{\alpha}[T]$ is a basic open set in $\prod_{\alpha}[T]$ which contains $f$ and which is disjoint from the image of $F_{\alpha}$. This shows that the image of $F_{\alpha}$ is closed in $\prod_{\alpha}[T]$. ∎
Corollary 2.15.
For every $\alpha<\omega_{1}$, $T^{1+\alpha}]$ admits a complete metric which makes it into a Polish space.
Доказательство.
This is because $F_{\alpha}$ embeds $T^{1+\alpha}]$ as a closed subspace of the Polish space $\prod_{1+\alpha}[T]$. ∎
@@ 3. Coding of the Borel Sets
Given a non-empty $\Pi_{\alpha+1}$-set $H=\bigcap_{n<\omega}H_{n}$, where each $H_{n}$ is a $\Sigma_{\alpha}$-set and given $c(H_{n})$ a set of branches of some subset $S_{n}$ of $T^{1+\alpha}$. We can MSO-define (see §4) such $S_{n}$ by taking all initial segments with respect to $T^{1+\alpha}$ of branches in $c(H_{n})$, or we could even take the whole $T^{1+\alpha}$, as we shall only use the branches that are actually in $c(H_{n})$ (we prefer calling the set $S_{n}$ for the ease of visualising the process, but an important point to retain for the future is that $S_{n}$ can be chosen in an MSO-definable way). We shall only proceed if the sequence $\langle H_{n}:\,n<\omega\rangle$ is $\subseteq$-decreasing, which we can do without loss generality since any $\Pi_{\alpha+1}$-set has a representation satisfying that property.
@@
For $\alpha=0$, in the case of closed or open sets, we have that $c(H)=H$ with the identification between $^{\omega}2$ and $2^{\omega}$ that is used throughout. We show how to determine if a set $H\subseteq[T]$ is closed. Let $s(H)=\{s\in T:\,(\exists\rho\in H)s\subseteq\rho\}$. If this set is a tree, then $H$ is closed and $s(H)=T_{H}$. To determine is $H$ is open, we determine if $2^{\omega}\setminus H$ is closed.
@@ 4.3. Translating into $S2S$
Coming back to our $\alpha$, we use these observations to give an MSO-representation of $(T^{\alpha},\leq_{\alpha})$ within $T$. By the induction hypothesis, we have an MSO-representation of each $(T^{\beta},\leq_{\beta})$ for $\beta<\alpha$ AND the elements of $T^{\beta}$ are represented by certain $\beta$-sequences of $T$. Let $(A_{\alpha},\prec_{\alpha})$ be an MSO-definable subset of $T$ ordered by $\prec_{\alpha}$ in the order type $\alpha$. We shall define a formula $G_{\alpha}(\varrho)$:
@@ 4.5. Subsets of $T^{1+\alpha}$ and branches in $T^{1+\alpha}]$
Let us first remark that for every $\alpha$ there is a formula ${\rm Path}_{1+\alpha}(\mathbf{B})$ such that for $B\subseteq T^{1+\alpha}$ we have ${\mathfrak{N}}_{2}\models(t{\rm Path}_{1+\alpha})(f{1+\alpha}(B))$ iff $B$ is a path of $T^{1+\alpha}$. This formula is
$(\forall x)(\forall y)(x\in\mathbf{B}\wedge y\in\mathbf{B}\!\implies\!x\leq_{\alpha}y\vee y\leq_{\alpha}x)].$

### v5

@@
I thank Joan Bagaria for the invitation to visit him at the University of Barcelona in April 2025. The conversation we had gave me a crucial insight. Namely, I showed Joan the idea of well-levelled orders and he proposed that this could probably be used in descriptive set theory. I went to work on this and connected it with Rabin’s work and solving the problem of MSO on the Borel sets, about which I had been thinking about on and off for a while. The idea of writing a joint paper on this did not converge, due to Joan’s other projects. I thank Andreas Blass, Ohad Drucker, Yair Hayut, Jouko Vддnдnen and Andrйs Villaveces on very useful comments on the preprint of the paper. In addition, I thank both the Set Theory Seminar and the Logic Seminar at the Hebrew University of Jerusalem for the active comments of the participants and for giving me an opportunity to present this work at its and mine alma mater.
@@ 2. Well-levelled orders and their basic properties
In the forward direction, let us write $P$ as the union of $\sim$-equivalence classes, which we shall enumerate as $\{C_{\alpha}:\alpha<\kappa\}$. To see that this is possible, apply the argument of (1) to conclude from ${\mathsf{wl}}(P)=0$ that every element of $P$ is above a unique minimal element and $\sim$-equivalent to it. Hence we can let $C_{\alpha}$ be the set of elements equivalent to a given minimal element, and then each $C_{\alpha}$ is a linear order, since $p\sim q$ implies that $p$ and $q$ are comparable. For $\alpha<\kappa$ let $q_{\alpha}$ be the unique minimal element of $C_{\alpha}$. Then we conclude as in (1) that $C_{\alpha}$ is the well-order $C_{\alpha}=\bigcup_{p\in C_{\alpha}}[q_{\alpha},p]$. We shall show that there is no relation between the elements of $C_{\alpha}$ and $C_{\beta}$ for $\alpha\neq\beta$.
@@
A path in $T^{\alpha}$ is a linearly ordered subset $\pi$ such that for no $\beta\leq\alpha$ does $\pi\upharpoonright\beta$ have an upper bound. Note that by the fact that $T^{\alpha}$ is well-founded, every path is order isomorphic to an ordinal. An $\omega$-path is a path that is order isomorphic to $\omega$. A branch is a path that is downward closed. (For $\alpha=1$ this coincides with the maximal paths, but not for $\alpha$ is general.) The set of all branches in $T^{\alpha}$ is denoted by $T^{\alpha}]$.
@@ 2.1. The Polish Topology of $T^{1+\alpha}]$
Suppose that $(X_{n},\psi_{n})_{n<\omega}$ is a sequence of non-empty complete metric spaces with metrics bounded by $1$. Then the product $\prod_{n<\omega}X_{n}$ admits a complete metric given by
$\rho(f,g)=\sum_{n<\omega}2^{-(n+1)}\psi_{n}(f(n),g(n)).$
@@
Theorem 2.13 can be adapted to show the following Lemma 2.14.111We thank Yair Hayut who during our talk on this paper at the Set Theory Seminar at the Hebrew University of Jerusalem on 14/01/2026 noticed a mistake in the previous version of the proof of Lemma 2.14. The mistake was due to the definition of a branch which did not fit what we have considered a branch in the rest of the paper.
Lemma 2.14.
For every $\alpha<\omega_{1}$, the space $T^{1+\alpha}]$ admits a complete compatible metric $d_{\alpha}$ bounded by 1.
@@
The proof is by induction on $\alpha$.
For $\alpha>0$, this is the classical fact that the Cantor set $[T]$ has a complete metric which generates the topology, while we have $T^{1}]=[T]$.
For $\alpha=\beta+1$, first note that if $\rho$ is a branch in $T^{1+\alpha}$, then $\rho\upharpoonright[1+\beta)$ is a branch in $T^{1+\alpha}$, by the definition of what a branch is.
We define $d_{\alpha}(f,g)=\frac{1}{2}\cdot[d_{\beta}(f\upharpoonright[1,\beta),g\upharpoonright[1,\beta))+d_{1}(f(\beta),g(\beta))]$. This function is clearly bounded by 1, non-negative and symmetric, and $d_{\alpha}(f,g)=0$ iff $f=g$. The triangle inequality follows by simple substitution and the use of the triangle inequalities for $d_{\beta}$ and $d_{1}$.
To see that the metric is complete, note that if $\langle f_{n}:\,n<\omega\rangle$ is a Cauchy sequence in $d_{\alpha}$, then $\langle f_{n}\upharpoonright[1,\beta):\,n<\omega\rangle$ is a Cauchy sequence in $d_{\beta}$, hence it converges to some $f\in[T^{1+\beta}]$. Similarly, $\langle f_{n}\upharpoonright(\beta):\,n<\omega\rangle$ is a Cauchy sequence in $d_{1}$, so it converges to some $g\in[T]$. Then $\langle f_{n}:\,n<\omega\rangle$ converges to the branch $\{t\!\frown\!s:\,t\in f\wedge s\in g\}$ in $T^{1+\alpha}]$.
For $\alpha>0$, let $\langle\beta_{n}:\,n<\omega\rangle$ be an increasing cofinal sequence in $\alpha$. Let $f,g\in[T^{\alpha}]$. Define $d_{\alpha}(f,g)$ using the formula similar to that of Theorem 2.13 with $\rho_{n}=d_{\beta_{n}}$, namely
$d_{\alpha}(f,g)=\sum_{n<\omega}2^{-(n+1)}d_{\beta_{n}}(f\upharpoonright[1,\beta_{n}),g\upharpoonright[1,\beta_{n})),$
but note that $f$ and $g$ here are not the elements of any product, but branches in $T^{\alpha}]$. The definition is justified since for every $f\in[T^{1+\alpha}]$ we have $f\upharpoonright[1+\beta_{n})$ is a branch in $T^{1+\beta_{n}}$.
The proof that this is a metric bounded by 1 is similar as in the successor case. To show that it is complete, suppose that $\langle f_{n}:\,n<\omega\rangle$ is a Cauchy sequence in $d_{\alpha}$. It follows that for every $\beta_{n}$, the sequence $\langle f_{n}\upharpoonright[1,\beta):\,n<\omega\rangle$ is a Cauchy sequence in $d_{\beta_{n}}$ and hence converges to some $\pi_{n}\in[T^{1+\beta_{n}}]$. Moreover, it follows by the definitions taken for $d_{\beta_{n}}$’s, that for $n\leq m$ we have $\pi_{n}=\{t\upharpoonright[1+\beta_{n}):\,t\in\pi_{m}\}$. Let
$\pi=\{t\in T^{\alpha}:\,(\forall n)t\upharpoonright[1+\beta_{n})\in\pi_{n}\}.$
We claim that $\pi$ is an element of $T^{\alpha}]$ and the limit of $\{\pi_{n}:\,n<\omega\}$. First let us show that it consists of pairwise compatible elements in $T^{\alpha}$, so suppose that $t,r\in\pi$ and let $k$ be the smallest such that $t\upharpoonright[\beta_{k},\alpha)=r\upharpoonright[\beta_{k},\alpha)\equiv\Lambda$. By the definition, $t\leq_{T^{\alpha}}r$ iff $t\upharpoonright[1+\beta_{k})\leq_{T^{1+\beta_{k}}}r\upharpoonright[1+\beta_{k})$ and vice versa. One of these possibilities must occur since $\pi_{k}$ is a branch in $T^{1+\beta_{k}}$. We similarly see that $\pi$ is downward closed. Let us calculate $d_{\alpha}(\pi,\pi_{n})$ for an arbitrary $n<\omega$. By the definition, this is $\leq\sum_{n\leq m<\omega}2^{-(m+1)}$, a quantity that converges to 0 as $n$ increases. Hence $\pi$ is the limit, as required to show. ∎
@@ 3. Coding of the Borel Sets
Given a non-empty $\Pi_{\alpha+1}$-set $H=\bigcap_{n<\omega}H_{n}$, where each $H_{n}$ is a $\Sigma_{\alpha}$-set and given $c(H_{n})$ a set of branches of some subset $S_{n}$ of $T^{1+\alpha}$. We can MSO-define (see §4) such $S_{n}$ by taking all initial segments with respect to $T^{1+\alpha}$ of branches in $c(H_{n})$, or we could even take the whole $T^{1+\alpha}$, as we shall only use the branches that are actually in $c(H_{n})$ (we prefer calling the set $S_{n}$ for the ease of visualising the process, but an important point to retain for the future is that $S_{n}$ can be chosen in an MSO-definable way. Note that from the inductive assumption that $c(H_{n})$ is MSO-definable, we can decide for any branch of $T^{1+\alpha}$ if it is in $c(H_{n})$ or not). We shall only proceed if the sequence $\langle H_{n}:\,n<\omega\rangle$ is $\subseteq$-decreasing, which we can do without loss generality since any $\Pi_{\alpha+1}$-set has a representation satisfying that property.
@@
For $\alpha=0$, in the case of closed or open sets, we have that $c(H)=H$ with the identification between $^{\omega}2$ and $2^{\omega}$ that is used throughout. We show how to determine if a set $H\subseteq[T]$ is closed. Let $s(H)=\{s\in T:\,(\exists\rho\in H)s\subseteq\rho\}$. This set is a tree and $H$ is closed iff $H=[s(H)]$, in which case $s(H)=T_{H}$. To determine is $H$ is open, we determine if $2^{\omega}\setminus H$ is closed.
@@ 4.3. Translating into $S2S$
Coming back to our $\alpha$, we use these observations to give an MSO-representation of $(T^{\alpha},\leq_{\alpha})$ within $T$. By the induction hypothesis, we have an MSO-representation of each $(T^{\beta},\leq_{\beta})$ for $\beta<\alpha$ AND the elements of $T^{\beta}$ are represented by certain $\beta$-sequences of the elements of $T$. Let $(A_{\alpha},\prec_{\alpha})$ be an MSO-definable subset of $T$ ordered by $\prec_{\alpha}$ in the order type $\alpha$. We shall define a formula $G_{\alpha}(\varrho)$:
@@ 4.5. Subsets of $T^{1+\alpha}$ and branches in $T^{1+\alpha}]$
Let us first remark that for every $\alpha$ there is a formula ${\rm Path}_{1+\alpha}(\mathbf{B})$ such that for $B\subseteq T^{1+\alpha}$ we have ${\mathfrak{N}}_{2}\models(t{\rm Path}_{1+\alpha})(f_{1+\alpha}(B))$ iff $B$ is a path of $T^{1+\alpha}$. This formula is
$\begin{split}[(\forall x)(\forall y)(x\in\mathbf{B}\wedge y\in\mathbf{B}\!\implies\!x\leq_{\alpha}y\vee y\leq_{\alpha}x)]\wedge\\ (\forall\beta<\alpha)\neg(\exists z)(\forall x)[x\in\mathbf{B}\!\implies\!x\upharpoonright\beta\leq z\upharpoonright\beta]].\end{split}$